OpenStudy (anonymous):

Can you help me see where I am stuck finding the zeros of this quadratic equation?

OpenStudy (anonymous):

\[3x^2 - 12x + 11 = 0\]

OpenStudy (anonymous):

quadratic formula for this one use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=3,b=-12,c=11\]

OpenStudy (anonymous):

My student solutions manual says I should obtain \[(6 \pm \sqrt{3}) \div 3\]

OpenStudy (anonymous):

indeed you should

OpenStudy (anonymous):

Let's walk through this, so I can see where I am tripping over my shoelaces.

OpenStudy (anonymous):

first you get \[\frac{-12\pm\sqrt{(-12)^2-4\times 3\times 11}}{6}\]

OpenStudy (anonymous):

Wait a sec. I see that I made an error in the first term under the radical. I will go back to the drawing board and re-attempt this one on my own.

OpenStudy (anonymous):

By the way, how are you producing the nice looking divisor line?

OpenStudy (anonymous):

\frac{a}{b} looks like \[\frac{a}{b}\]

OpenStudy (anonymous):

Got it. I made simple mistakes in my substitutions, then got hung up on trying to do algebra with unduly complicated figures.

OpenStudy (anonymous):

THANKS @satellite73, you got me unstuck!